OpenStudy (anonymous):

find derivative: integral x=3 to sinx : 1/2 e^(t^2)dt

OpenStudy (amistre64):

find derivative of an integral?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

\[\frac{d}{dx}\int_{a(x)}^{b(x)}f(t)dt\to f(b)b'-f(a)a'\]

OpenStudy (anonymous):

find derivative of the function which is an integral

OpenStudy (amistre64):

\[\int_{a}^{b}f(t)dt=F(b)-F(a)\] \[\frac{d}{dx}\ F(b)-F(a)=f(b)b'-f(a)a'\]

OpenStudy (anonymous):